Sorry -4(x- 11)-34=10-12 could you show how you solved this?

Answer by Alan3354(69443)

-4(x- 11)-34=10-12
-------------
-4x + 44 - 34 = -2
-4x + 10 = -2
Subtract 10
-4x = -12
4x = 12
x = 3